# Symmetry-aware spectral bounds for the electronic structure Hamiltonian

### v1

@@ Abstract
where $\hat{S}_{+}=\sum_{p}\hat{a}^{\dagger}_{p\alpha}\hat{a}_{p\beta}$ and $\hat{S}_{-}=(\hat{S}_{+})^{\dagger}$. This Hamiltonian also remains invariant to any transformation involving a well-behaved function of these symmetries, $f(\hat{N},\hat{S}_{z},\hat{S}^{2})$. It is important to note that while the Numerical Results Section of this manuscript only considers particle number and spin symmetries, specific problem instances might also possess additional symmetries that should be included. For instance, permutational and point-group symmetries for specific molecular structures, such as rotation, reflection, and inversion, should also be considered. In addition, periodic symmetries in the context of crystalline systems could also greatly reduce the $\ell_{1}$ norm [36, 37]. As a result, these bounds suggest that the extensive scaling behavior of block-invariant $\ell_{1}$ norms should significantly differ from what has been observed in previous numerical studies.
@@
Scaling analysis. While symmetry-aware spectral bounds provide insights into the performance limits of quantum algorithms based on Hamiltonian oracle models, they do not provide any insight into the extensive scaling behavior with respect to system size. Previous numerical analysis of the scaling behavior of the $\ell_{1}$ norm for the electronic structure Hamiltonian has been performed based on sparse, double factorization, and THC [22, 38] representations. It has been shown to scale polynomially, $\lambda=\mathcal{O}(\text{poly}(N_{\mathrm{orb}}))$, defined with respect to the number of orbitals, $N_{\mathrm{orb}}$. The symmetry-aware spectral bounds presented in this manuscript suggest a different scaling that should incorporate some dependence with respect to the total particle number $\eta$ or other symmetry-sector eigenvalue. In Appendix C, we present theoretical results for the scaling behavior of the symmetry-aware spectral bounds for the electronic structure Hamiltonian, highlighted by the upper bound dependence, $\Delta_{\mu}\leq\mathcal{O}(g(N_{\mathrm{orb}})\eta^{2})$. Our results find that an underlying spatial orbital dependence originates from the function $g(N_{\mathrm{orb}})$, which is heavily dependent on the type of basis set. In the numerical results section below, we show that the family of correlation-consistent basis sets demonstrate a sublinear dependence, $g(N_{\mathrm{orb}})=\mathcal{O}(N_{\mathrm{orb}}^{x})$ with $x\leq 1$, for different chemical systems. The following section outlines a numerically tractable procedure for computing these bounds for large systems and presents quantitative results for several benchmark cases.
@@
Numerical Results. To numerically determine the three spectral bounds ($\Delta_{\mu}/2,\Delta_{\mathrm{s}}/2,$ and $\Delta/2$) for molecular systems, the smallest and largest eigenvalues, denoted by $E^{\mu}_{\mathrm{min}}$ and $E^{\mu}_{\mathrm{max}}$, must be computed in every symmetry sector $\mu$. In the context of electronic structure theory, we consider the three symmetries, $(\hat{N},\hat{S}^{2},\hat{S}_{z})$, outlined above with each sector uniquely defined by the triplet, $\mu=(\eta,s,m_{s})$, where $\eta$ equals the total number of electrons, while $s$ and $m_{s}$ represent the total spin and its projection, respectively. In the absence of external magnetic fields and spin-orbit coupling, as considered throughout this manuscript, the quantum numbers $\eta$ and $s$ are sufficient to uniquely identify all energy eigenvalues. While an ideal calculation would provide these eigenvalues with full configuration interaction (FCI)-level precision, the computational demand becomes prohibitive for large system sizes. To address this, we propose a numerically tractable method that involves performing orbital optimization to identify the determinant that minimizes or maximizes the energy within each sector, similar to a restricted Hartree-Fock level calculation suitable in both closed-shell and open-shell cases. It is important to emphasize that this method does not provide multi-reference capabilities and, therefore, may not accurately describe certain physical states, such as open-shell singlets (detailed in Appendix D). Despite this limitation, our method adheres to the variational principle, ensuring that the estimated spectral ranges for each sector satisfy the inequality,
@@
This guarantees that our numerical estimates serve as a conservative lower bound to the actual FCI spectral range. While more expensive methodologies (e.g., CISD, CCSD, or DMRG) could be used to provide more accurate assessments of the spectral bounds, we found that the orbital-optimized methodology remains useful in providing non-trivial estimates of the spectral range as well as its scaling behavior (see Appendix D for details). It is worth noting that for large system sizes, the calculation of the spectral bound ($\Delta/2$) and symmetry bound ($\Delta_{s}/2$) becomes computationally intensive due to the total number of symmetry sectors, scaling as $\mathcal{O}(N^{2})$.
@@
[…]
@@
In addition to providing the three spectral bounds ($\Delta_{\mu}/2,\Delta_{\mathrm{s}}/2,$ and $\Delta/2$) for the total Hamiltonian in the last column, we have also computed the spectral bound for the 1-body and 2-body terms separately (the details of the decomposition is outlined in Appendix C). The second last column (in gray) denotes the sum of the 1-body and 2-body bounds, which we call the incoherent spectral bound. This column should serve as the reference point for most contemporary tensor factorization techniques, including double factorization and tensor hypercontraction (THC) [21, 22], due to the fact that these methods factorize the 1-body and 2-body terms independently (see Appendix C).
@@
As highlighted in Table 1, the discrepancy between the incoherent and coherent spectral bounds suggests that applying tensor compression methods to a dressed four-index tensor, which integrates both 1-body and 2-body contributions, might provide further reduction in the $\ell_{1}$ norm. Our prior research demonstrated the efficacy of this concept within the context of observable estimation [41], achieving up to a four-fold reduction in the $\ell_{1}$ norm for specific cases. This decrease in the $\ell_{1}$ norm results from the natural cancellation of energy contributions within the dressed tensor. Nonetheless, a thorough examination of this approach is warranted, particularly in the context of estimating ground-state energies, with careful consideration of its impact on $C_{\mathcal{W}[H]}$.
@@
Traditional tensor compression methods, like double factorization and THC, which do not incorporate symmetry information, are constrained by the spectral bound ($\Delta/2$) found in the top row. As a point of reference, the published $\ell_{1}$ norm values based on THC correspond to 306 (Reiher et al. [39]) and 1202 (Lee et al. [22]), in units of Hartree. These numbers suggest that additional enhancements to THC, such as applying regularization techniques [33, 25], refining initial starting guesses, or relaxing the constraints on the THC rank, could at most achieve improvement factors of around 2.0 and 1.6, respectively. On the other hand, the symmetry-based bounds suggest that future tensor compression techniques that exploit symmetry information might be able to achieve an approximate improvement factor of up to 6 and 21 for the FeMoco-based benchmark systems outlined above.
[…]
@@
The coherent spectral bound is lower than its incoherent counterpart in every instance. This disparity becomes increasingly pronounced with larger basis sets, exhibiting nearly an order of magnitude difference between the symmetry-sector bounds. Consistent with the theoretical findings from the previous section, our analysis reveals markedly different scaling behaviors between the regular bound, symmetry bound, and the symmetry-sector bound, as evidenced by their distinct slopes. This leads to a nearly three-order-of-magnitude difference between the regular spectral bound ($\Delta/2$) and the symmetry-sector bound ($\Delta_{\mu}/2$) in the cc-pV5Z basis. In particular, the symmetry-sector bound $\Delta_{\mu}/2$ exhibits a sublinear slope, $g(N)=\mathcal{O}(N^{x})$, with $x\leq 1$ across all three molecular systems.
@@
Thermodynamic Limit Scaling. To conclude our numerical analysis, we also compute the spectral bounds of a linear hydrogen chain at half-filling as a function of the number of hydrogen atoms, $N_{H}$, ranging from 10 to 50. The base geometry consists of a linear array of hydrogen atoms separated by 0.74 Ångstrom expanded in the minimal STO-6G basis set. In Fig. 2, we find that while there is nearly an order of magnitude difference between the regular spectral bound ($\Delta/2$) and the symmetry bound ($\Delta_{\mathrm{s}}/2$), the associated difference between the symmetry bound and the symmetry sector bound ($\Delta_{\mu}/2$) is found to be nearly negligible. We attribute this to the half-filling sector typically having the largest spectral range among all symmetry sectors for this benchmark system. Since the gap between conventional and symmetry-aware bounds is more pronounced in the low-filling regime (as also highlighted in Fig. 1), symmetry-shifted tensor compression techniques should be sufficient to saturate the achievable $\ell_{1}$ norm bounds in the half-filling regime.
@@
Discussion. This study introduces symmetry-aware spectral bounds relevant to a wide variety of quantum algorithms in physics and chemistry. For qubitization-based QPE, we find an asymptotic Toffoli complexity of $\mathcal{O}(\Delta_{\mu}N)$, assuming a linear $N$ scaling of the qubitization walk operator as observed in THC [22]. Combined with the upper bound scaling of the symmetry-sector bound, $\Delta_{\mu}=\mathcal{\tilde{O}}(N^{x}\eta^{2})$, derived in this work, we predict a final Toffoli complexity of $\mathcal{\tilde{O}}(N^{1+x}\eta^{2})$. Unlike previous work, this scaling has an explicit dependence on the number of particles $\eta$ making it much more favourable in the continuum and complete basis set limit. In contrast, first-quantization methods using plane waves have a Toffoli complexity of $\mathcal{O}(\eta^{8/3}N^{1/3})$ [42]. This raises the question of whether block-encoding strategies with explicit knowledge of symmetries could be constructed. Such dependence would enhance the competitiveness of second-quantized methods using localized Gaussian orbitals compared with first-quantization techniques in these limits.

### v2

@@ III.1 Electronic Structure Hamiltonian
where $\hat{S}_{+}=\sum_{p}\hat{a}^{\dagger}_{p\alpha}\hat{a}_{p\beta}$ and $\hat{S}_{-}=(\hat{S}_{+})^{\dagger}$. This Hamiltonian also remains invariant to any transformation involving a well-behaved function of these symmetries, $f(\hat{N},\hat{S}_{z},\hat{S}^{2})$. It is important to note that while the Numerical Results Section of this manuscript only considers particle number and spin symmetries, specific problem instances might also possess additional symmetries that should be included. For instance, permutational and point-group symmetries for specific molecular structures, such as rotation, reflection, and inversion, should also be considered. In addition, periodic symmetries in the context of crystalline systems could also greatly reduce the 1-norm [36, 37]. As a result, these bounds suggest that the extensive scaling behavior of block-invariant 1-norms should significantly differ from what has been observed in previous numerical studies.
@@
Scaling analysis. While symmetry-aware spectral bounds provide insights into the performance limits of quantum algorithms based on Hamiltonian oracle models, they do not provide any insight into the extensive scaling behavior with respect to system size. Previous numerical analysis of the scaling behavior of the 1-norm for the electronic structure Hamiltonian has been performed based on sparse, double factorization, and THC [22, 38] representations. It has been shown to scale polynomially, $\lambda=\mathcal{O}(\text{poly}(N_{\mathrm{orb}}))$, defined with respect to the number of orbitals, $N_{\mathrm{orb}}$. The symmetry-aware spectral bounds presented in this manuscript suggest a different scaling that should incorporate some dependence with respect to the total particle number $\eta$ or other symmetry-sector eigenvalue. In Appendix A, we present theoretical results for the scaling behavior of the symmetry-aware spectral bounds for the electronic structure Hamiltonian, highlighted by the upper bound dependence, $\Delta_{\mu}\leq\mathcal{O}(g(N_{\mathrm{orb}})\eta^{2})$. Our results find that an underlying spatial orbital dependence originates from the function $g(N_{\mathrm{orb}})$, which is heavily dependent on the type of basis set. In the numerical results section below, we show that the family of correlation-consistent basis sets demonstrate a sublinear dependence, $g(N_{\mathrm{orb}})=\mathcal{O}(N_{\mathrm{orb}}^{x})$ with $x\leq 1$, for different chemical systems. The following section outlines a numerically tractable procedure for computing these bounds for large systems and presents quantitative results for several benchmark cases.
IV Numerical evaluation of symmetry-aware spectral bounds
@@
To numerically determine the three spectral bounds ($\Delta_{\mu}/2,\Delta_{\mathrm{s}}/2,$ and $\Delta/2$) for molecular systems, the smallest and largest eigenvalues, denoted by $E^{\mu}_{\mathrm{min}}$ and $E^{\mu}_{\mathrm{max}}$, must be computed in every symmetry sector $\mu$. In the context of electronic structure theory, we consider the three symmetries, $(\hat{N},\hat{S}^{2},\hat{S}_{z})$, outlined above with each sector uniquely defined by the triplet, $\mu=(\eta,S,m_{s})$, where $\eta$ equals the total number of electrons, while $S$ and $m_{s}$ represent the total spin and its projection, respectively. In the absence of external magnetic fields and spin-orbit coupling, as considered throughout this manuscript, the quantum numbers $\eta$ and $S$ are sufficient to uniquely identify all energy eigenvalues. While an ideal calculation would provide these eigenvalues with full configuration interaction (FCI)-level precision, the computational demand becomes prohibitive for large system sizes. To address this, we propose a numerically tractable method that involves performing orbital optimization to identify the determinant that minimizes or maximizes the energy within each sector, similar to a restricted Hartree-Fock level calculation suitable in both closed-shell and open-shell cases. It is important to emphasize that this method does not provide multi-reference capabilities and, therefore, may not accurately describe certain physical states, such as open-shell singlets (detailed in Appendix B). Despite this limitation, our method adheres to the variational principle, ensuring that the estimated spectral ranges for each sector remains a lower bound to the exact FCI solution (see Figure 1),
@@
This guarantees that our numerical estimates serve as a conservative lower bound to the actual FCI spectral range. While more expensive methodologies (e.g., CISD, CCSD, or DMRG) could be used to provide more accurate assessments of the spectral bounds, we found that the orbital-optimized methodology remains useful in providing non-trivial estimates of the spectral range as well as its scaling behavior (see Appendix B for details). It is worth noting that for large system sizes, the calculation of the spectral bound ($\Delta/2$) and symmetry bound ($\Delta_{s}/2$) becomes computationally intensive due to the total number of symmetry sector. In a fixed basis set with $N_{\mathrm{orb}}$ spatial orbitals, the total number of electrons will be limited to the range $[0,2N_{\mathrm{orb}}]$ across the entire Hilbert space. In the absence of external magnetic fields or spin-orbit coupling, the latter resulting from relativistic adjustments to the Hamiltonian, the energy eigenvalues characterized by constant $\eta$ and $S$, but varying $m_{s}$, will be degenerate. This implies that the total number of symmetry sectors $(\eta,S)$ with unique energy eigenvalues will be given by,
$\frac{(N_{\mathrm{orb}}+1)(N_{\mathrm{orb}}+2)}{2}.$ (24)
While this number is only quadratic, it renders the numerical evaluation of $\Delta/2$ and $\Delta_{s}/2$ quite resource-intensive for large basis sets.
IV.1 Orbital Optimization Procedure
$E_{\mathrm{HF}}(\boldsymbol{\kappa})\!=\!\sum_{i}^{\eta_{\alpha}}h_{ii}(\boldsymbol{\kappa})+\sum_{i}^{\eta_{\beta}}h_{ii}(\boldsymbol{\kappa})+\!\frac{1}{2}\!\!\left(\!\sum_{i,j}^{\eta_{\alpha},\eta_{\alpha}}\!\!g_{iijj}(\boldsymbol{\kappa})+\!\sum_{i,j}^{\eta_{\beta},\eta_{\beta}}\!\!g_{iijj}(\boldsymbol{\kappa})+2\!\!\sum_{i,j}^{\eta_{\alpha},\eta_{\beta}}\!\!g_{iijj}(\boldsymbol{\kappa})-\!\sum_{i,j}^{\eta_{\alpha},\eta_{\alpha}}\!\!g_{ijji}(\boldsymbol{\kappa})-\!\sum_{i,j}^{\eta_{\beta},\eta_{\beta}}\!\!g_{ijji}(\boldsymbol{\kappa})\right)\!\!,$ (25)
with orbital-transformed one-electron $h_{ij}$ and two-electron $g_{ijkl}$ tensor coefficients,
$h_{ij}(\boldsymbol{\kappa})$ $=\sum_{pq}h_{pq}[\exp(-\boldsymbol{\kappa})]_{pi}[\exp(-\boldsymbol{\kappa})]_{qj},$ (26)
$g_{ijkl}(\boldsymbol{\kappa})$ $=\sum_{pqrs}g_{pqrs}[\exp(-\boldsymbol{\kappa})]_{pi}[\exp(-\boldsymbol{\kappa})]_{qj}[\exp(-\boldsymbol{\kappa})]_{rk}[\exp(-\boldsymbol{\kappa})]_{sl},$ (27)
defined with respect to the unitary matrix, $\exp(-\boldsymbol{\kappa})$, written in terms of the anti-Hermitian matrix, $\boldsymbol{\kappa}$, where $\boldsymbol{\kappa}^{\dagger}=-\boldsymbol{\kappa}$. To improve the stability of the optimization procedure (especially relevant for finding the maximun eigenvalue), we used the pre-orthogonalized one-electron $h_{pq}$ and two-electron integrals $g_{pqrs}$,
$h_{pq}$ $=\sum_{\mu\nu}h_{\mu\nu}[S^{-\frac{1}{2}}]_{\mu p}[S^{-\frac{1}{2}}]_{\nu q},$ (28)
$g_{pqrs}$ $=\sum_{\mu\nu\rho\sigma}g_{\mu\nu\rho\sigma}[S^{-\frac{1}{2}}]_{\mu p}[S^{-\frac{1}{2}}]_{\nu q}[S^{-\frac{1}{2}}]_{\rho r}[S^{-\frac{1}{2}}]_{\sigma s},$ (29)
as our initial starting guess. These quantities are defined with respect to the atomic orbital overlap matrix ($S_{\mu\nu}$) as well as the one-electron ($h_{\mu\nu}$) and two-electron ($g_{\mu\nu\rho\sigma}$) integrals defined in the atomic orbital basis. For the optimization of $N_{\mathrm{orb}}$ real spatial orbitals, it is sufficient to consider the real part of $\boldsymbol{\kappa}$, which contains only $N_{\mathrm{orb}}(N_{\mathrm{orb}}-1)/2$ parameters. The maximum eigenvalue, $E_{\mathrm{max}}^{(\eta_{\alpha},\eta_{\beta})}$, can be found with the same numerical minimizer by taking the negative of the one-electron and two-electron integrals. It is important to note that for RHF and ROHF techniques, the associated wave function will be an eigenfunction of the total spin operator, $\hat{S}^{2}$, with an expectation value equal to,
$\braket{\hat{S}^{2}}=\left(\frac{\eta_{\alpha}-\eta_{\beta}}{2}\right)\!\!\left(\frac{\eta_{\alpha}-\eta_{\beta}}{2}+1\right),$ (30)
assuming $\eta_{\alpha}\geq\eta_{\beta}$. As a result, the spin quantum number $s$ is equal to the spin-projection quantum number $|m_{s}|$. Although this might appear restrictive, this implies that in scenarios devoid of magnetic fields and spin-orbit interactions, adjusting $\eta_{\alpha}$ and $\eta_{\beta}$ suffices to encompass all possible symmetry sectors characterized by unique energy eigenvalues.
V Numerical Results
@@
     BeH2
     H2O
     NH3
Figure 2: The spectral bound $\Delta/2$, symmetry bound $\Delta_{s}/2$ and symmetry-sector bound $\Delta_{\mu}/2$ as a function of spatial orbitals, $N_{\mathrm{orb}}$, with correlation-consistent basis sets (cc-pVDZ, cc-pVTZ, cc-pVQZ, cc-pV5Z) for three molecules BeH2, H2O, and NH3, with all values reported in units of Hartree.
@@
In addition to providing the three spectral bounds ($\Delta_{\mu}/2,\Delta_{\mathrm{s}}/2,$ and $\Delta/2$) for the total Hamiltonian in the last column, we have also computed the spectral bound for the 1-body and 2-body terms separately (the details of the decomposition is outlined in Appendix A). The second last column (in gray) denotes the sum of the 1-body and 2-body bounds, which we call the incoherent spectral bound. This column should serve as the reference point for most contemporary tensor factorization techniques, including double factorization and tensor hypercontraction (THC), due to the fact that these methods factorize the 1-body and 2-body terms independently [21, 22].
@@
As highlighted in Table 1, the discrepancy between the incoherent and coherent spectral bounds suggests that applying tensor compression methods to a dressed four-index tensor, which integrates both 1-body and 2-body contributions, might provide a further reduction in the 1-norm. Our prior research demonstrated the efficacy of this concept within the context of observable estimation [41], achieving up to a four-fold reduction in the 1-norm for specific cases. This decrease in the 1-norm results from the natural cancellation of energy contributions within the dressed tensor. Nonetheless, a thorough examination of this approach is warranted, particularly in the context of estimating ground-state energies, with careful consideration of its impact on $C_{\mathcal{W}[H]}$.
@@
Traditional tensor compression methods, like double factorization and THC, which do not incorporate symmetry information, are constrained by the spectral bound ($\Delta/2$) found in the top row. As a point of reference, the published 1-norm values for FeMoco based on THC correspond to 306 (Reiher et al. [39]) and 1202 (Lee et al. [22]) and for the P450 Hamiltonian of 388.9 (Goings et al. [33]), in units of Hartree. These numbers suggest that additional enhancements to THC, such as applying regularization techniques [33, 25], refining initial starting guesses, or relaxing the constraints on the THC rank, could at most achieve improvement factors of around 2.0 and 1.6, respectively. On the other hand, the symmetry-based bounds suggest that tensor compression techniques that exploit symmetry information in the future might be able to achieve an approximate improvement factor of up to 6 and 21 for the FeMoco-based benchmark systems outlined above. A recently introduced symmetry-shifted version of double factorization (SCDF) was shown to violate the conventional bound ($\Delta/2$) and achieved a 1-norm of 78.0 for FeMoco (Reiher) and 111.3 for P450 (Rocca et al. [rocca2024reducing]), but remained the symmetry-shifted spectral bound ($\Delta_{\mathrm{s}}/2$) as expected.
Complete Basis Set Limit Scaling. To further highlight the differences between these bounds, we compute the scaling behavior of the three spectral bounds ($\Delta_{\mu}/2,\Delta_{\mathrm{s}}/2,$ and $\Delta/2$) for BeH2, H2O, and NH3. These molecules and their respective geometries were selected based on prior research [34, 35] because they are small enough to enable complete basis set extrapolation analysis. The numerical analysis is conducted as a function of basis set size within the correlation-consistent (cc-pVnZ) basis sets, ranging from cc-pVDZ to cc-pV5Z. The size of each basis set is quantified explicitly by the number of spatial orbitals, denoted as $N_{\mathrm{orb}}$. For each of these systems, the symmetry sector bound ($\Delta_{\mu}/2$) is fixed to the singlet ($S=0$) sector with the total electron counts of $6$, $10$, and $10$, respectively. The resulting data are illustrated in Fig. 2, where both the incoherent (orange markers) and regular (blue markers) spectral bounds are presented for comparative analysis.
@@
The coherent spectral bound is lower than its incoherent counterpart in every instance. This disparity becomes increasingly pronounced with larger basis sets, exhibiting nearly an order of magnitude difference between the symmetry-sector bounds. Consistent with the theoretical findings from the previous section, our analysis reveals markedly different scaling behaviors between the regular bound, the symmetry bound, and the symmetry-sector bound, as evidenced by their distinct slopes. This leads to a nearly three-order-of-magnitude difference between the regular spectral bound ($\Delta/2$) and the symmetry-sector bound ($\Delta_{\mu}/2$) in the cc-pV5Z basis. In particular, the symmetry-sector bound $\Delta_{\mu}/2$ exhibits a sublinear slope, $g(N)=\mathcal{O}(N^{x})$, with $x\leq 1$ across all three molecular systems.
@@
Thermodynamic Limit Scaling. To conclude our numerical analysis, we also compute the spectral bounds of a linear hydrogen chain at half-filling as a function of the number of hydrogen atoms, $N_{H}$, ranging from 10 to 50. The base geometry consists of a linear array of hydrogen atoms separated by 0.74 Ångstrom expanded in the minimal STO-6G basis set. In Fig. 3, we find that while there is nearly an order of magnitude difference between the regular spectral bound ($\Delta/2$) and the symmetry bound ($\Delta_{\mathrm{s}}/2$), the associated difference between the symmetry bound and the symmetry sector bound ($\Delta_{\mu}/2$) is found to be nearly negligible. We attribute this to the half-filling sector typically having the largest spectral range among all symmetry sectors for this benchmark system. Since the gap between conventional and symmetry-aware bounds is more pronounced in the low-filling regime (as also highlighted in Fig. 2), symmetry-shifted tensor compression techniques should be sufficient to saturate the achievable 1-norm bounds in the half-filling regime.
VI Conclusion
@@
This study introduces symmetry-aware spectral bounds relevant to a wide variety of quantum algorithms in physics and chemistry. For qubitization-based QPE, we find an asymptotic Toffoli complexity of $\mathcal{O}(\Delta_{\mu}N_{\mathrm{orb}})$, assuming a linear $N_{\mathrm{orb}}$ scaling of the qubitization walk operator as observed in THC [22]. Combined with the upper bound scaling of the symmetry-sector bound, $\Delta_{\mu}=\mathcal{\tilde{O}}(N_{\mathrm{orb}}^{x}\eta^{2})$, derived in this work, we predict a final Toffoli complexity of $\mathcal{\tilde{O}}(N_{\mathrm{orb}}^{1+x}\eta^{2})$. Unlike previous work, this scaling has an explicit dependence on the number of particles $\eta$ making it much more favorable in the continuum and complete basis set limit. In contrast, first-quantization methods using plane waves have a Toffoli complexity of $\mathcal{O}(\eta^{8/3}N_{\mathrm{orb}}^{1/3})$ [42]. This raises the question of whether block-encoding strategies with explicit knowledge of symmetries could be constructed. Such dependence would enhance the competitiveness of second-quantized methods using localized Gaussian orbitals compared with first-quantization techniques in these limits.